THEORETICAL MODELING OF INFRARED SPECTRA OF ASPIRIN AND ITS DEUTERATED DERIVA-TIVE

M. BOCZAR, \underline{M.J. WÓJCIK}, K. SZCZEPONEK, D. JAMRÓZ, A. ZIȨBA, AND BOZENA KAWALEK, Faculty of Chemistry, Jagiellonian University, 30-060 Kraków, Ingardena 3, Poland.

Theoretical simulation of the ν_{s} stretching band is presented for aspirin (acetylsalicylic acid) and its OD derivative at 300 and 77 K. The simulation takes into account an adiabatic coupling between the high-frequency O-H(D) stretching and the low-frequency intermolecular O...O stretching modes, linear and quadratic distortions of the potential energy for the low-frequency vibrations in the excited state of the O-H(D) stretching vibration, resonance interaction between two hydrogen bonds in the dimer, and Fermi resonance between the O-H(D) stretching and the overtone of the O-H(D) bending vibrations. The effect of deuteration and the temperature has been successfully reproduced by our model calculations.Infrared, far-infarared, Raman and low-frequency Raman spectra of the polycrystalline aspirin have been measured. The geometry and experimental frequencies is compared with the results of our B3LYP/6-31++G** calculations.